### v1

# Low-Rank Continual Personalization of Diffusion Models

Abstract
Recent personalization methods for diffusion models, such as Dreambooth, allow fine-tuning pre-trained models to generate new concepts. However, applying these techniques across multiple tasks in order to include, e.g., several new objects or styles, leads to mutual interference between their adapters. While recent studies attempt to mitigate this issue by combining trained adapters across tasks after fine-tuning, we adopt a more rigorous regime and investigate the personalization of large diffusion models under a continual learning scenario, where such interference leads to catastrophic forgetting of previous knowledge. To that end, we evaluate the naïve continual fine-tuning of customized models and compare this approach with three methods for consecutive adapters’ training: sequentially merging new adapters, merging orthogonally initialized adapters, and updating only relevant parameters according to the task. In our experiments, we show that the proposed approaches mitigate forgetting when compared to the naïve approach.
11footnotetext: Equal contribution.
1 Introduction
Deep Diffusion Probabilistic Models (DDPMs) [28] have revolutionized image generation, with state-of-the-art performance in synthesizing highly detailed objects and blending them with complex styles. Moreover, personalization techniques such as Dreambooth [24], combined with the Low-Rank Adaptation (LoRA) [12], introduce the possibility to easily customize large-scale pre-trained models. In those techniques, a selected subset of model parameters is updated with just a few training examples to allow for the generation of objects not present in the original training set.
Unfortunately, continuous customization of a generative model over several tasks leads to the catastrophic forgetting [6] of previously encoded knowledge due to mutual interference between tasks’ adapters. Recent studies try to mitigate this issue through additional regularization [27] or by careful merging of the trained adapters [26, 7]. For such merging to succeed, all adapters have to be available at the same time to properly scale [26] or select [18] the resulting models’ weights. In practice, with a bigger number of tasks, this assumption is often unfeasible, as storing numerous copies of model weights, even in the form of a sparse low-rank adaptation matrix, is memory-consuming (e.g. approximately 8 LoRA weight matrices with rank $64$ are equal to the size of all adapted parameters in Stable Diffusion XL’s [21] model). Therefore, in this work, we propose to study the effectiveness of merging techniques under the strict continual learning regime where only the model or model with a single adapter are passed between tasks.
In a series of experiments, we evaluate (1) Naïve continual fine-tuning of the low-rank adapter with three merging methods mitigating forgetting of concepts and styles: (2) consecutive merging of task-specific adapters, (3) merging LoRAs initialized with orthogonal weights, and (4) merging through a selection of weights with the highest magnitude for the task [18]. Our experiments indicate that adding multiple adapters in a Naïve way can lead to a situation where a model converges towards its base form, while all the evaluated techniques mitigate this issue.
2 Related work
Diffusion models personalization and customization
The topic of Diffusion Models (DMs) personalization has been broadly studied in recent years. Several works, including Textual Inversion [8] learn new word embeddings to represent custom and unique concepts. This idea is further extended by several approaches [17, 16], with additional regularization [33], or features mapping [32]. On the other hand, DreamBooth [24] fine-tunes the entire model to incorporate specific subjects into generated images. This idea is further extended in HyperDreamBooth [25] where a hypernetwork [10] is used to generate personalized weights. Recent personalization techniques employ Low-Rank Adaptation (LoRA) [12] to reduce the computational requirements by fine-tuning small, low-rank matrices and keeping the base model frozen. Nevertheless, simply using LoRA does not prevent interference. Therefore, ZipLoRA [26] optimizes separately trained adapters for style and object to simplify their merging, and B-LoRA [7] targets different attention layers of diffusion models for both style and object adapters. On top of those, weights2weights [5] introduces a personalized adapter space, enabling the creation of entirely new personalized models. OFT [22] and Orthogonal Adaptation [20] leverage orthogonalization to merge models seamlessly without compromising fidelity.
Generating style with diffusion models
Apart from introducing new objects to the model, there is a line of research aiming for style generation and editing in DDPMs. StyleDrop [29] leverages the Muse transformer to fine-tune text-to-image models for generating style-specific images. StyleAligned [11] employs shared attention to maintain a consistent style across generated images, while Style Similarity [30] provides a framework for style retrieval from diffusion models’ training data and introduces CSD, an image encoder that accurately captures the style of images. UnlearnCanvas [36] offers a comprehensive dataset for evaluating machine unlearning in stylized image generation. For knowledge localization, Diff-QuickFix [2] and LOCOEDIT [1] enable precise editing of diffusion models’ knowledge of specific styles by identifying style-relevant layers.
Continual Learning of Diffusion Models
There are works using DMs in CL scenarios. In [35], the authors explore if contemporary CL methods prevent forgetting in diffusion models. In [9], DDPM is used as a source of rehearsal examples, while [4] extends this idea with classifier guidance towards easy-to-forget examples. Similarly, in [15], authors use an externally trained text-to-image diffusion model for the same purpose. To mitigate artifacts introduced with generative replay, especially with low-rank adaptation, in [27], the authors introduce C-LoRA, where consecutive tasks are learned with a self-regularization technique that penalizes updates corresponding to already altered weights.
Model merging
The concept of editing models through arithmetic operations applied on models’ weights is explored in [13], where authors describe mathematical operations between task vectors, defined as the difference between the weights of the fine-tuned and the base model. To limit the interference between tasks, recent methods replace straightforward averaging with summation of trimmed parameters [34], Fisher-information-based merging [19] or with guidance from discriminative model [14]. In this work, we adapt the MagMax method [18] that is used to continually train the CLIP model by selecting, from several task vectors, weights with the highest magnitude.
3 Methods
We evaluate four different approaches to continual model personalization. Notation-wise, we denote $W_{0}$ as the base model weights, while $(A_{t},B_{t})$ represent the fine-tuned low-rank adapter resulting from personalizing DM on the $t$-th task. New adapter weights $(A_{t},B_{t})$ are calculated through fine-tuning on data from task $t$, given the frozen base model weights $W_{0}$, and an adapter fine-tuned over $t-1$ tasks denoted as $(A_{1..t-1},B_{1..t-1})$. The final model can be calculated as $W_{1...T}=W_{0}+B_{1...T}A_{1...T}$. All the evaluated strategies differ in initialization or merging strategy, which we overview in Fig. 1.
Figure 1: Comparison of initialization and merging in four evaluated methods. The expressions in the adapters refer to how they are initialized with a new task.
(1) Naïve continual fine-tuning of LoRA adapters.
In this naïve approach, we sequentially fine-tune the adapter weights from the previous task: $A_{t-1}$, $B_{t-1}$ without reinitializing them. The intermediate adapters are not merged into the base model.
(2) Continual adapters merging and standard LoRA reinitialization.
In the second approach, we create a new LoRA adapter for each task with a standard initialization technique ($A\sim N(0,I),B=0$). Then, at the end of each task, we merge the adapter with the base model and pass it to the next task.
(3) Continual adapters merging and orthogonalized reinitialization.
We further extend the merging and reinitialization approach and limit the interference between tasks by pre-directing the fine-tuning process toward non-conflicting adapters by initializing $B_{t}$ weights with zeros and setting $A_{t}$ weights as orthogonal to $A_{1..t-1}$ using Singular Value Decomposition (SVD). We decompose the $i$-th column of $A_{1..t-1}$ as:
${A^{(i)}_{1..t-1}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}$ (1)
and select the last row of the conjugate transpose of the right singular vectors $\mathbf{V}$, corresponding to the smallest singular value.
(4) Magnitude-based selection of LoRA weights
In the last approach, we adapt the magnitude-based parameter selection introduced in [18]. To that end, after each task, we select parameters with the highest magnitude when comparing already merged adapters $(A^{\prime}_{1..t-1},B^{\prime}_{1..t-1})$ with the current one $(A_{t},B_{t})$. The resulting merged adapter is passed together with a base model to the new task. Task-related adapters are always initialized in the standard way, while for the frozen Base model, we use a version with parameters updated by the merged adapter.
4 Experiments
Experimental setup
In our experiments, we use the DreamBooth [24] dataset for fine-tuning on objects, and UnlearnCanvas [36] for styles. We perform fine-tuning on the $10$ consecutive tasks and average results over 4 random tasks ordering, with 2 random seeds each. To measure the performance of the model, we use cosine similarity to the reference image with DINO [3] and CLIP [23] feature extractors for object alignment measure and DINO and CSD [31] for style similarity. We report the performance of the final model fine-tuned over all 10 tasks with the Average Score and Average Forgetting metrics. More details about the experimental setup are provided in the Appendix A.
Continual Personalization Methods Comparison
In Tab.1 and Tab.2, we present the results for the continual object and style personalization, respectively. For both tasks, we observe that the merge & initialization method, either with or without orthogonalization, outperforms naïve continual fine-tuning. On the other hand, magnitude-based merging has the lowest average forgetting at the expense of lower general performance as a result of the lower adaptation capabilities of the method. While magnitude-based weights selection yields state-of-the-art performance in discriminative tasks [18], selecting a limited number of weights hinders the precise merging of new generative traits. Detailed analysis in Appendix B on performance after each task shows that the merge & orthogonalization approach exhibits the best plasticity at the expense of the highest forgetting.
Average Score $\overline{S}_{T}\,(\uparrow)$ Average Forgetting $\overline{F}_{T}\,(\downarrow)$
Adapter fine-tuning method CLIP-I DINO CLIP-I DINO
Base model (reference) $0.586$ $0.304$
Naïve Continual Fine-Tuning $0.670_{\pm.013}$ $0.402_{\pm.029}$ $0.063_{\pm.013}$ $0.144_{\pm.039}$
Merge & Initialization $\boldsymbol{0.675}_{\pm.005}$ $\boldsymbol{0.457}_{\pm.011}$ $\underline{0.026}_{\pm.003}$ $\underline{0.056}_{\pm.009}$
Merge & Orthogonal Initialization $\underline{0.673}_{\pm.014}$ ${0.403}_{\pm.025}$ $0.072_{\pm.015}$ $0.162_{\pm.033}$
Magnitude-based Merging $0.643_{\pm.002}$ $\underline{0.408}_{\pm.006}$ $\boldsymbol{0.018}_{\pm.002}$ $\boldsymbol{0.036}_{\pm.006}$
Table 1: Average Score and Average Forgetting in CLIP-I and DINO object alignment metrics for our continual personalization methods. The best results are in bold, while the second best are underlined.
Average Score $\overline{S}_{T}\,(\uparrow)$ Average Forgetting $\overline{F}_{T}\,(\downarrow)$
Adapter fine-tuning method CSD DINO CSD DINO
Base model (reference) $0.088$ $0.146$
Naïve Continual Fine-Tuning $0.345_{\pm.032}$ $0.249_{\pm.010}$ $0.184_{\pm.039}$ $0.136_{\pm.013}$
Merge & Initialization $\boldsymbol{0.385}_{\pm.019}$ $\boldsymbol{0.285}_{\pm.019}$ $\underline{0.131}_{\pm.019}$ $\underline{0.085}_{\pm.021}$
Merge & Orthogonal Initialization $\underline{0.349}_{\pm.014}$ $\underline{0.252}_{\pm.012}$ $0.204_{\pm.013}$ $0.141_{\pm.013}$
Magnitude-based Merging $0.289_{\pm.015}$ $0.240_{\pm.009}$ $\boldsymbol{0.093}_{\pm.013}$ $\boldsymbol{0.052}_{\pm.015}$
Table 2: Average Score and Average Forgetting in CSD and DINO style alignment metrics for our continual personalization methods. The best results are in bold, while the second best are underlined.
To further analyze the results, in Fig 2 (left), we present a comparison of different methods’ performance on the first task after continual adaptation on the given number of tasks. We can observe that the naïve approach decays to the performance of the baseline model, while other techniques can mitigate forgetting better. In Fig 2 (right), we present the possible reason for adapters’ degradation.
We measure the extent of the mutual interference between tasks during training. To that end, we create a vector build of directions that oppose the adapters of a task after training on subsequent tasks and plot the ratio of this vector’s norm to the original task vector norm. An explanation of calculating the degradation of adapters, along with the detailed results of our evaluation for the style adaptation task, are presented in Appendix B.
Refer to caption Refer to caption
Figure 2: DINO score on the first task over continual fine-tuning on next tasks for the object customization (left) and the ratio of opposed parameters vector norm to the task adapter norm (right).
5 Conclusions
In this work, we analyze the continual adaptation capabilities of a DreamBooth method coupled with the Low-Rank Adaptation (LoRA) technique. We evaluate three different approaches to merging and initialization of the individual adapters. We show that naïve continual training of a LoRA leads to catastrophic forgetting, while other techniques can mitigate this issue, which originates from conflicts in adapters that are sequentially trained.
6 Acknowledgements
This research was supported by NCN OPUS grant no. 2023/51/B/ST6/03004. The computing resources were provided by PL-Grid Infrastructure grant no. PLG/2024/017176.
Appendix A Experimental setup
A.1 Datasets
DreamBooth To fine-tune the model on personalized objects, the DreamBooth [24] dataset is used, which contains images of 30 objects (9 of them being dogs and cats, and 21 being other objects) belonging to 15 different classes. Each object is represented by several images (from 4 up to 6), which represent it in either a different environment or from a different camera angle. Due to these quantity limits and dataset set-up, we use the same dataset of reference images for both evaluation and fine-tuning.
UnlearnCanvas For style-related tasks, we use UnlearnCanvas collection [36], which contains 20 objects (each with 20 images) in 60 different styles. When evaluating methods for the style tasks, we select $10$ images representing $10$ different objects in the same style but the ones that were not used during training.
We generated datasets for our experiments by randomly selecting classes and objects. For object-related tasks, we selected 10 classes, each represented by a unique object and paired with a unique token. For style-based tasks, we selected 10 distinct styles, each paired with one object and five corresponding images. Object names, classes, and styles were uniquely selected to avoid repetition. Selected classes are presented in Table 3.
Style name byzantine, palette knife, ink art, blossom season, artist sketch,
winter, dadaism, dapple, meteor shower, pastel
Object name sneaker, boot, backpack, stuffed animal, candle, can, glasses, bowl, cartoon, teapot
Table 3: Objects and styles names used during experiments.
A.2 Finetuning
We perform the experiments in a class-incremental setup, as illustrated in Figure 1. Starting with a base text-to-image Stable Diffusion XL model, we sequentially train it on the following tasks. After each task, depending on the approach, we either merge the obtained LoRA weights with the base model and pass it to the next task or store only the LoRA weights to initialize adapters for the subsequent steps.
We train the model on $T=10$ tasks. For each object task, we use distinct DreamBooth tokens, while for styles, we retain the real names. The learning rate is set to 1e-5, with 850 train steps and a batch size of 1. During fine-tuning, we don’t leverage the prior preservation regularization [24].
In the original paper introducing LoRA [12], the authors investigate how both the rank and the target matrices of the transformer’s self-attention module, to which the LoRA is applied, influence method efficiency. They show that, while the impact of rank $r$ is small, it is recommended to apply the LoRA to $W^{Q}$, $W^{K}$, $W^{V}$, and $W^{O}$. In our experiments, we similarly apply LoRA to these targets within the self-attention weights of the U-Net and set $r=64$.
All four methods for low-rank continual personalization methods were evaluated with two seeds (00, $5$) and four task ordering seeds (00, $5$, $10$, $42$).
A.3 Evaluation metrics
To evaluate fine-tuned model generations, we use image alignment metrics to calculate the cosine similarity between embeddings of generated and reference images. For subject alignment, we use features extracted from the DINO [3] and CLIP [23] models. When assessing style, we continue using DINO but replace CLIP with CSD [31], as CLIP has demonstrated a bias towards image content, making it unsuitable for style-related tasks. When sampling from the text-to-image diffusion model, we generate $N=8$ images from each of $P=5$ prompts and calculate the mean cosine similarity of models’ embeddings between each of $N\cdot P=40$ samples and the object/style reference images. When generating an image, we do it with $50$ denoising diffusion steps.
When investigating model performance in a continual learning regime, we use two metrics: average score $\overline{S}_{T}=\frac{1}{T}\sum_{j=1}^{T}S_{j}^{T}$ (with $S_{j}^{i}$ being performance on $j$-th task of a model that was trained on $i$ consecutive tasks) and average forgetting $\overline{F}_{T}=\frac{1}{T-1}\sum_{j=1}^{T-1}\max_{1\leq k\leq T}(S_{j}^{k}-S_{j}^{T})$, with $S_{j}^{T}$ being score (CLIP-I, DINO, CSD) on $j$-th task of a model that was trained on $T$ consecutive tasks.
Appendix B Additional results
For style personalization, we also observe in Fig 3 (right) the origin of each method forgetting, which is the magnitude of conflicts introduced to adapters by fine-tuning on the next tasks. We see that both methods that achieve the lowest forgetting in Fig. 3 (left) - magnitude-based merging and merge & initialization - exhibit the least increasing ratio of vector norms over time. Note that for the $9$-th task, we observe plot degradation that comes from calculating conflicts according to adapters that were trained from the base model without including adapters from tasks trained later.
Refer to caption Refer to caption
Figure 3: Performance on the first task (left) and the ratio of conflicting vector norm to introduced adapter parameter (right) over continual fine-tuning on next tasks for the style personalization.
The model generally achieves its highest performance on each task immediately after fine-tuning, but this performance tends to degrade as new tasks are added. Despite this decline, our evaluated techniques can retain knowledge from previous tasks. As presented in Fig. 67 for object personalization and Fig. 89 for style personalization, we can observe that the merge & orthogonal initialization method exhibits the highest plasticity in all metrics for both objects and styles. For some cases (tasks 1 and 2 in merge & initialization method in Fig. 6), we can observe backward transfer, a situation where the performance on the previous task improves thanks to the adaptation to the new task.
For measuring $t$-th task adapter’s degradation with the next tasks incoming, we leverage two vectors of parameters: (1) task vector representing parameters introduced to the model only by fine-tuning it on that one task - calculated by subtraction of $t-1$-th adapter from the $t$-th adapter - and (2) next $m$ tasks’ vector denoting parameters introduced to the model starting from $t+1$-th task till the $t+m$-th task - calculated by subtraction of $t+m$-th adapter from the $t$-th adapter - and zero its parameters that keep the same sign as the corresponding ones in task vector. Finally, we determine the task degradation by the ratio of the norm of (2) the conflicted next tasks’ vector to the norm of (1) the task vector. We average the ratios over all the adapters, not only the first one.
Appendix C Prompts for styles
Due to the complex nature of style personalization, this task is characterized by the easy overwriting of previous styles when new ones come adapted by the Dreambooth method. This fact is extremely harmful especially in the context of continued learning, where it implies catastrophic forgetting of the model. In our experiments, we note that a significant problem in the context of personalization of styles is the way they are expressed in the form of a prompt. To this end, we conduct how different forms of prompts affect the average score, average forgetting, and task scores of a diffusion model learning next styles sequentially.
We evaluate models using six different prompts while training them across five consecutive style-related tasks, aiming to identify the prompt that achieved the highest accuracy and lowest forgetting. The results are summarized in Table 4. We also measure the average CSD alignment score and its changes across successive tasks - results in Table 5. We observe that using classic Dreambooth prompts like ’image in the $S^{*}$ style’ introduces worse results than specifying the full names of the styles, especially with the token ’style’ inside.
Prompt Template DINO $\overline{S}_{T}\,(\uparrow)$ DINO $\overline{F}_{T}\,(\downarrow)$ CSD $\overline{S}_{T}\,(\uparrow)$ CSD $\overline{F}_{T}\,(\downarrow)$
image of {} in {token} style 0.188 0.215 0.237 0.272
image of {} in {name} style 0.289 0.249 0.418 0.359
image of {} in {token} 0.260 0.279 0.377 0.433
image of {} in {name} 0.292 0.237 0.423 0.348
{token} image of {} 0.268 0.249 0.387 0.389
{name} image of {} 0.299 0.201 0.446 0.275
Table 4: Performance metrics average score and average forgetting across different style prompts templates for DINO and CSD.
Prompt Template $\mathbf{Task_{1}}$ $\mathbf{Task_{2}}$ $\mathbf{Task_{3}}$ $\mathbf{Task_{4}}$ $\mathbf{Task_{5}}$
image of {} in {token} style $0.365$ $0.267$ $0.213$ $0.117$ $0.084$
image of {} in {name} style 0.503 $0.362$ $0.321$ $0.186$ $0.261$
image of {} in {token} $0.491$ $0.303$ $0.236$ $0.115$ $0.168$
image of {} in {name} $0.493$ $0.371$ $0.337$ $0.191$ $0.261$
{token} image of {} $0.487$ $0.350$ $0.250$ $0.130$ $0.231$
{name} image of {} $0.487$ 0.410 0.397 0.255 0.289
Table 5: The average value of the CSD metric demonstrates how models perform on the selected tasks, where 1 represents the most recently trained task, and 5 represents the earliest task.
Appendix D Example samples
Figure 4 presents object images generated by the models fine-tuned on all 10 consecutive tasks using our continual learning methods, while Figure 5 presents it for styles.
Figure 4: Images generated by the model fine-tuned on 10 tasks. The first left column represents the subsequent tasks, each represented by an image illustrating the object to learn. We show how each method sequentially performs using the prompt templates: ’a $V^{*}$ wearing a santa hat’ and ’a $V^{*}$ in a jail’.
Figure 5: Images generated by the model fine-tuned on 10 tasks. The first left column represents the subsequent tasks, each represented by an image illustrating the style to learn. We show how each method sequentially performs using the prompt templates: ’{style_name} image of a wizard’ and ’{style_name} image of a mirror’.
(a) Naïve continual fine-tuning of LoRA adapters
(b) Continual adapters merging and standard LoRA reinitialization
(c) Continual adapters merging and orthogonalized reinitialization
(d) Magnitude-based selection of LoRA weights
Figure 6: Heatmap of CLIP-I alignment scores for each task in continual object personalization.
(a) Naïve continual fine-tuning of LoRA adapters
(b) Continual adapters merging and standard LoRA reinitialization
(c) Continual adapters merging and orthogonalized reinitialization
(d) Magnitude-based selection of LoRA weights
Figure 7: Heatmap of DINO alignment scores for each task in continual object personalization.
(a) Naïve continual fine-tuning of LoRA adapters
(b) Continual adapters merging and standard LoRA reinitialization
(c) Continual adapters merging and orthogonalized reinitialization
(d) Magnitude-based selection of LoRA weights
Figure 8: Heatmap of DINO alignment scores for each task in continual style personalization.
(a) Naïve continual fine-tuning of LoRA adapters
(b) Continual adapters merging and standard LoRA reinitialization
(c) Continual adapters merging and orthogonalized reinitialization
(d) Magnitude-based selection of LoRA weights
Figure 9: Heatmap of CSD alignment scores for each task in continual style personalization.